A washer-dryer combination costs $1200. If the washer costs $220 more than the dryer, what does each appliance cost separately?
Answer by faceoff57(108)

w = 220 + d
d + (220 + d) = 1200
d + 220 + d = 1200
2d + 220 = 1200
2d = 1200 - 220
2d = 980
d = 490
the dryer is $490.
the washer is $710.